Optical action page:

· Open actions
· Finished actions
   
· Optics Main
· Mechanical Main

 

Contacts:

· Maarten Blanken mfb@ing.iac.es

 

This page will show you the action of the optics group.

Optical action page


If you feel that I have forgotten any optical issue in here, please send an email to: mfb@ing.iac.es

Open actions up
Date Action Fault number Status Comment
17/02/2003 CO2 cleaning WHT, INT, JKT      
03/02/2003 Reflectivity measurement WHT, INT, JKT      
20/01/2003 CO2 cleaning WHT, INT, JKT      
06/01/2003 Reflectivity measurement WHT, INT, JKT      
23/12/2002
CO2 cleaning INT, JKT



09/12/2002
Reflectivity measurement WHT, INT, JKT



16/09/2002
Filter measurement

17/09/2002 Checked the spectrograph at SLO.

It is not working correct. Varian is contacted, waiting for reply.
20/09/2002 contacted Varian for support.

20/11/2002 Ordered new lamp from varian

Measure the TAURUS filters to be used in the WHT PFC
12/06/2002
Update Filter database

12/06/2002 Did not update filter database. Waiting till new version

is out. Contacts Robert Greimel and Craige Bevil.


12/06/2002
Filter check

12/06/2002 Checked filters and cleaned when necesary.


Finished actions up
Date Action Fault number Status Comment Date when finished
15-17/12/2002
WHT aluminising, Nasmyth washing

Done
Reflectivity result page

15-17/12/2002
23/11/2002 WHT, INT, JKT CO2 cleaning
 
Done
Reflectivity result page

Could not take pictures of WHT because of GRACE building

28/11/2002
04/11/2002
 
Done
Reflectivity result page
14/11/2002
28/10/2002
WHT, INT CO2 cleaning

Done
Reflectivity result page
30/10/2002
24/10/2002
JKT washing

Done
Reflectivity result page
24/10/2002
07/10/2002
Reflectivity measurement WHT, INT, JKT

Done
Reflectivity result page
08/10/2002
02/10/2002
INGRID alignment check with Sue Worswick
PIMS issue 121

Found a detector tilt of 400 microns over the y direction and 100 microns over the x directrion. Also have to look at the PS positions.
08/10/2002
23/09/2002 CO2 cleaning WHT, INT, JKT

Reflectivity result page 25/09/2002
16/09/2002
Optical stock WHT cabinet

Done
Need optical tissue is WHT cabinet
25/09/2002
13/09/2002
Filter check INT WFC
#15304
13/09/2002 checked filter orientation. A few filters are wrongly orientated.
25/09/2002 Re orientated filters #199, #207, #210, #220
13/09/2002 Ask Danny Lennon if orientation fix does not change science output.
25/09/2002
13/09/2002
Reflectivity measurement WHT, INT, JKT


Reflectivity result page
13/09/2002
13/08/2002


Clean corrector plate of RoboDIMM.

Done washed it with water

25/09/2002
 


ING Logo

Last Updated: 09/01/2007
Tibor Agócs tibor@ing.iac.es